OpenStudy (anonymous):

Find the value of x to the nearest tenth. Problem: tan2 = 4 over x

OpenStudy (igreen):

Is that \(\sf tan(2^o) = \dfrac{4}{x}\)

OpenStudy (anonymous):

Yes.

OpenStudy (igreen):

Okay, plug in tan 2 degrees into your calculator, what do you get?

OpenStudy (anonymous):

0.0349207695

OpenStudy (igreen):

Yes, so we have: \(\sf 0.0349207695 = \dfrac{4}{x}\) Multiply 'x' to both sides: \(\sf 0.0349207695x = 4\) Divide 0.0349207695 to both sides.

OpenStudy (igreen):

4 / 0.0349207695 = ?

OpenStudy (anonymous):

Gives me the answer of 114.5450131046. Ahhh, I get it now, thanks. c:

OpenStudy (igreen):

Yep, you got it.

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

Thanks. :)